You are obtaining a $250,000 FRM at a 8% annual percentage rate. What is the monthly payment?


amortized over 30 years
Answer #1

Monthly payment = [principal amount * ( r / n ) ] / [ 1 - (1 + r/n )(- n * amortized year)]

Principal amount= 250000

r = 0.08

Amortized over = 30 year

Monthly payment = [ 250000 * (0.08/ 12) ] / [1- (1+ 0.08/12)(-12*30)

  = ( 250000 * 0.0066) / [ 1- (1.0066)(-12*30) ]

= 1650 / (1 -0.0936)

= 1650 / 0.9064

= 1820.388